# Effects of Flagellar Morphology on Swimming Performance and Directional Control in Microswimmers

### v1

@@ III.2 Hydrodynamic Interaction between Cell Body and Flagellum
Figures 6(a) and (c) present the forward speed of the microswimmer in the circumference-pitch plane, where Fig. 6(a) represents the forward speed when neglecting the hydrodynamic interaction between the cell body and the flagellum, while Fig. 6(c) includes this interaction. Similarly, Figs. 6(b) and (d) illustrate the swimming efficiency of the microswimmer under the same conditions. In particular, when the hydrodynamic interaction between the cell body and the flagellum is considered, the forward speed of the microswimmer decreases, but the swimming efficiency increases. The opposing rotational directions of the cell body and the flagellum result in an ambient flow field where $\Omega_{x}^{\infty}<0$. According to the relationship $\mathcal{R}_{xx}^{rr}(\Omega_{x}-\Omega_{x}^{\infty})=T_{x}$, the torque exerted on the cell body by the motor increases when this hydrodynamic interaction is taken into account. As shown in Figs. 5(b) and (e), the ambient translational flow field generated by the flagellum also satisfies $U_{x}^{\infty}<0$. Thus, based on the relationship $\mathcal{R}_{xx}^{tt}(U_{x}-U_{x}^{\infty})=F_{x}$, the propulsive force exerted by the flagellum on the microswimmer increases when considering this hydrodynamic interaction. In this context, the propulsive force with hydrodynamic interaction is approximately $1.5$ times greater than that when this interaction is neglected. According to the swimming efficiency formula Eq. 7, the variation in swimming efficiency results from the competition between the propulsive force and the torque.
@@ III.3 Optimal Morphology of Flagella
Firstly, we assess the impact of flagellar contour length on the forward speed and swimming efficiency of the microswimmer. Using a helix radius of $R=0.2$ $\mu\mathrm{m}$, we present the contours of forward speed and swimming efficiency for a freely swimming microswimmer in Fig. 7. The analysis reveals that, for varying contour lengths, there is an optimal pitch angle, approximately $\pi/5$, which maximizes either forward speed or swimming efficiency. Additionally, the optimal pitch angles for contour lengths of $\Lambda=2$ $\mu\mathrm{m}$ and $\Lambda=4$ $\mu\mathrm{m}$ are slightly higher than those for the other contour lengths.
@@
We examine various helix radii, $R$, to calculate the maximum forward speed and maximum swimming efficiency across different contour lengths and pitch angles, along with their corresponding optimal pitch angles, as illustrated in Fig. 8. As shown in Fig. 8(a), the maximum forward speed increases with both the helix radius and contour length. However, once the helix radius exceeds $0.3$ $\mu\mathrm{m}$, the maximum efficiency reaches a plateau and does not improve further, as depicted in Fig. 8(b). Therefore, for a cell body radius of $R_{b}=1$ $\mu\mathrm{m}$, the optimal helix radius is approximately in the range of $0.2\leq R\leq 0.3$ $\mu\mathrm{m}$, and the corresponding optimal pitch angles range from $35^{\circ}$ to $45^{\circ}$. Moreover, for the majority of flagellar contour lengths, it is observed that the optimal pitch angle increases with the helix radius, as shown in Figs. 8(c) and (d).
@@
Microswimmers need to capture nutrients to survive in fluid environments Guasto, Rusconi, and Stocker (2012); Yawata et al. (2014). As shown in Figs. 2, and 4, the trajectories of microswimmers exhibit a circular helix, allowing them to navigate by adjusting the orientation of their flagella. In addition to considering forward speed and swimming efficiency, it is crucial to evaluate their directional capability for survival in these environments. To quantitatively characterize this capability, we introduce the yaw angle $\beta$, defined as the angle between the flagellar axis and the direction of forward locomotion. Fig. 9(a) demonstrates that a smaller helix radius is associated with a smaller yaw angle, indicating enhanced directionality. Furthermore, Fig. 8 shows that as the helix radius decreases, both forward speed and swimming efficiency decline. Analyzing Fig. 9(a) further confirms that a smaller flagellum helix radius correlates with stronger directional control. Additionally, Fig. 9(b) indicates that when $R<0.3$ $\mu\mathrm{m}$, the helical diameter of the trajectory is less than $26$ $\mathrm{n}\mathrm{m}$, which is comparable to the size of the nutrients necessary for the microswimmer’s survival. By integrating factors such as forward speed, swimming efficiency, yaw angle, and diameter of the helical trajectory, we conclude that the optimal range for the flagellar helix radius is $0.2\leq R\leq 0.3$ $\mu\mathrm{m}$.
@@
Figure 11 demonstrates the existence of an optimal pitch angle for the flagellum that maximizes either the forward speed or swimming efficiency of the microswimmer. We present contours of forward speed and swimming efficiency with respect to various flagellar contour lengths, from which we extract the maximum forward speed, maximum swimming efficiency, and their corresponding optimal pitch angles, as shown in Fig. 11. From Fig. 11(a), it is clear that the filament radius has a minimal impact on the maximum forward speed of the microswimmer. However, maximum swimming efficiency significantly decreases with smaller filament radii, as illustrated in Fig. 11(b). This suggests that while a smaller filament radius can enhance swimming efficiency, when the contour length of the flagellum exceeds $9$ $\mu\mathrm{m}$, the improvement in swimming efficiency becomes negligible. Therefore, longer flagellar contour lengths do not necessarily yield better performance for the microswimmer.
When the filament radius is less than $0.07$ $\mu\mathrm{m}$, the pitch angle for maximum forward speed is slightly lower than that for maximum swimming efficiency, with a difference not exceeding $4.5^{\circ}$. This minor variation allows the microswimmer to make slight adjustments to the pitch angle, facilitating a transition between flagellar morphologies optimized for either maximum forward speed or maximum swimming efficiency. This adaptability is essential for enhancing the survival capability of the microswimmer. This also explains why the Young’s modulus of most flagella is relatively high, a low Young’s modulus would be detrimental to bacterial survival Darnton and Berg (2007); Son, Guasto, and Stocker (2013); Vogel and Stark (2013). From Figs. 11(c) and (d), it is evident that when the filament radius falls within the range $0.01\leqslant a\leqslant 0.02$ $\mu\mathrm{m}$, the pitch angles corresponding to both maximum forward speed and optimal swimming efficiency are approximately $35^{\circ}\leqslant\theta\leqslant 40^{\circ}$. This pitch angle range aligns with the typical morphology of the flagellum Koyasu and Shirakihara (1984); Fujii, Shibata, and Aizawa (2008); Spagnolie and Lauga (2011).
@@ IV Summary and Conclusions
During swimming, the flagella generate a flow field that opposes the forward motion of the microswimmer, while the cell body produces a flow field in the same direction as the forward movement. Such hydrodynamic interaction between the cell body and the flagella is of great importance. Our findings indicate that considering this interaction significantly increases the propulsive force generated by the flagella. Although this interaction may reduce the forward speed, it improves swimming efficiency.
The morphology of flagella is essential for determining the forward velocity, swimming efficiency, and directional stability of microswimmers. Research indicates that an optimal flagellar morphology exists that maximizes either forward velocity or swimming efficiency. The optimal contour length of the flagellum is around $10$ $\mu\mathrm{m}$. Beyond this length, additional increases in contour length do not significantly enhance the swimming efficiency of the microswimmer. Although the radius of the filaments has a relatively minor effect on forward velocity, reducing the filament radius significantly enhances swimming efficiency. The optimal helix radius for microswimmers is approximately $0.2$ to $0.3$ $\mu\mathrm{m}$. Beyond this range, increasing the helix radius does not further improve swimming efficiency; rather, it increases the yaw angle and the diameter of the helical trajectory, adversely affecting directional stability. At a helix radius of $R=0.2$ $\mu\mathrm{m}$, the optimal pitch angle for maximum forward speed is slightly lower, by no more than $4.5^{\circ}$, than that for maximum swimming efficiency. This flexibility enables microswimmers to transition between maximum forward speed and maximum swimming efficiency by fine-tuning the pitch angle, with the optimal pitch angle range being approximately $30^{\circ}$ to $45^{\circ}$. Therefore, the flagellum only needs to make minor adjustments to its morphology to adapt to different living environments, which is why its Young’s modulus is relatively high.

### v2

@@ III.2 Hydrodynamic Interaction between Cell Body and Flagellum
Figures 6(a) and (c) present the forward speed of the microswimmer in the circumference-pitch plane, where Fig. 6(a) represents the forward speed while neglecting the hydrodynamic interaction between the cell body and the flagellum, while Fig. 6(c) includes this interaction. Similarly, Figs. 6(b) and (d) illustrate the swimming efficiency of the microswimmer under the same conditions. In particular, when the hydrodynamic interaction between the cell body and the flagellum is considered, the forward speed of the microswimmer decreases, but the swimming efficiency increases. The opposing rotational directions of the cell body and the flagellum result in an ambient flow field where $\Omega_{x}^{\infty}<0$. According to the relationship $\mathcal{R}_{xx}^{rr}(\Omega_{x}-\Omega_{x}^{\infty})=T_{x}$, the torque exerted on the cell body by the motor increases when this hydrodynamic interaction is taken into account. As shown in Figs. 5(b) and (e), the ambient translational flow field generated by the flagellum also satisfies $U_{x}^{\infty}<0$. Thus, based on the relationship $\mathcal{R}_{xx}^{tt}(U_{x}-U_{x}^{\infty})=F_{x}$, the propulsive force exerted by the flagellum on the microswimmer increases when considering this hydrodynamic interaction. In this context, the propulsive force with hydrodynamic interaction is approximately $1.5$ times greater than that when this interaction is neglected. According to the swimming efficiency formula Eq. 7, the variation in swimming efficiency results from the competition between the propulsive force and the torque.
@@ III.3 Optimal Morphology of Flagella
Firstly, we assess the impact of flagellar contour length on the forward speed and swimming efficiency of the microswimmer. Using a helix radius of $R=0.2$ $\mu\mathrm{m}$, we present the contours of forward speed and swimming efficiency for a freely swimming microswimmer in Fig. 7. The analysis reveals that for varying contour lengths, there is an optimal pitch angle, approximately $\pi/5$, which maximizes either forward speed or swimming efficiency. Additionally, the optimal pitch angles for contour lengths of $\Lambda=2$ $\mu\mathrm{m}$ and $\Lambda=4$ $\mu\mathrm{m}$ are slightly higher than those for the other contour lengths.
@@
We examine various helix radii, $R$, to calculate the maximum forward speed and maximum swimming efficiency across different contour lengths and pitch angles, along with their corresponding optimal pitch angles, as illustrated in Fig. 8. As shown in Fig. 8(a), the maximum forward speed increases with both the helix radius and contour length. However, once the helix radius exceeds $0.3$ $\mu\mathrm{m}$, the maximum efficiency reaches a plateau and does not improve further, as depicted in Fig. 8(b). Therefore, for a cell body radius of $R_{b}=1$ $\mu\mathrm{m}$, the optimal helix radius is approximately in the range of $0.2\leq R\leq 0.3$ $\mu\mathrm{m}$, and the corresponding optimal pitch angles range from $30^{\circ}$ to $45^{\circ}$. Moreover, for the majority of flagellar contour lengths, it is observed that the optimal pitch angle increases with the helix radius, as shown in Figs. 8(c) and (d).
@@
Microswimmers need to capture nutrients to survive in fluid environments Guasto, Rusconi, and Stocker (2012); Yawata et al. (2014). As shown in Figs. 2 and 4, the trajectories of microswimmers exhibit a circular helix, allowing them to navigate by adjusting the orientation of their flagella. In addition to considering forward speed and swimming efficiency, it is crucial to evaluate their directional capabilities for survival in these environments. To quantitatively characterize this capability, we introduce the yaw angle $\beta$, defined as the angle between the flagellar axis and the direction of forward locomotion. Fig. 9(a) demonstrates that a smaller helix radius is associated with a smaller yaw angle, indicating enhanced directionality. Furthermore, Fig. 8 shows that when the helix radius $R\leq 3$ $\mu\mathrm{m}$, both forward speed and swimming efficiency decrease as the helix radius decreases. Analyzing Fig. 9(a) further confirms that a smaller flagellum helix radius correlates with stronger directional control. Additionally, Fig. 9(b) indicates that when $R<0.3$ $\mu\mathrm{m}$, the helical diameter of the trajectory is less than $26$ $\mathrm{n}\mathrm{m}$, which is comparable to the size of the nutrients necessary for the microswimmer’s survival. As the helical radius of the microswimmer’s trajectory decreases, its efficiency in capturing nutrients increases. By integrating factors such as forward speed, swimming efficiency, yaw angle, and diameter of the helical trajectory, we conclude that the optimal range for the flagellar helix radius is $0.2\leq R\leq 0.3$ $\mu\mathrm{m}$.
@@
Figure 11 demonstrates the existence of an optimal pitch angle for the flagellum that maximizes either the forward speed or swimming efficiency of the microswimmer. We present contours of forward speed and swimming efficiency with respect to various flagellar contour lengths, from which we extract the maximum forward speed, maximum swimming efficiency, and their corresponding optimal pitch angles, as shown in Fig. 11. From Fig. 11(a), it is clear that the filament radius has a minimal impact on the maximum forward speed of the microswimmer. However, maximum swimming efficiency significantly decreases as the filament radii increase, as illustrated in Fig. 11(b). This suggests that while a smaller filament radius can enhance swimming efficiency, when the contour length of the flagellum exceeds $9$ $\mu\mathrm{m}$, the improvement in swimming efficiency becomes negligible. Therefore, longer flagellar contour lengths do not necessarily yield better performance for the microswimmer.
When the filament radius is less than $0.07$ $\mu\mathrm{m}$, the pitch angle for maximum forward speed is slightly lower than that for maximum swimming efficiency, with a difference not exceeding $4.5^{\circ}$. This minor variation allows the microswimmer to make slight adjustments to the pitch angle, facilitating a transition between flagellar morphologies optimized for either maximum forward speed or maximum swimming efficiency. This adaptability is essential for enhancing the survival capability of the microswimmer. This also explains why the Young’s modulus of most flagella is relatively high, as a low Young’s modulus would be detrimental to bacterial survival Darnton and Berg (2007); Son, Guasto, and Stocker (2013); Vogel and Stark (2013). From Figs. 11(c) and (d), it is evident that when the filament radius falls within the range $0.01\leqslant a\leqslant 0.02$ $\mu\mathrm{m}$, the pitch angles corresponding to both maximum forward speed and optimal swimming efficiency are approximately $30^{\circ}\leqslant\theta\leqslant 45^{\circ}$. This pitch angle range aligns with the typical morphology of the flagellum Koyasu and Shirakihara (1984); Fujii, Shibata, and Aizawa (2008); Spagnolie and Lauga (2011).
@@ IV Summary and Conclusions
During swimming, the flagella generate a flow field that opposes the forward motion of the microswimmer, while the cell body produces a flow field in the same direction as the forward movement. Such hydrodynamic interaction between the cell body and the flagella is crucial. Our findings indicate that considering this interaction significantly increases the propulsive force generated by the flagella. Although this interaction may reduce the forward speed, it enhances swimming efficiency.
The morphology of flagella is essential for determining the forward velocity, swimming efficiency, and directional stability of microswimmers. Research indicates that an optimal flagellar morphology exists that maximizes either forward velocity or swimming efficiency. The optimal contour length of the flagellum is around $10$ $\mu\mathrm{m}$. Beyond this length, additional increases in contour length do not significantly enhance the swimming efficiency of the microswimmer. Although the radius of the filaments has a relatively minor effect on forward velocity, reducing the filament radius significantly enhances swimming efficiency. The optimal helix radius for microswimmers is approximately $0.2$ to $0.3$ $\mu\mathrm{m}$. Beyond this range, increasing the helix radius does not further improve swimming efficiency; rather, it increases the yaw angle and the diameter of the helical trajectory, adversely affecting the directional stability. At a helix radius of $R=0.2$ $\mu\mathrm{m}$, the optimal pitch angle for maximum forward speed is slightly lower, by no more than $4.5^{\circ}$, than that for maximum swimming efficiency. This flexibility enables microswimmers to transition between maximum forward speed and maximum swimming efficiency by fine-tuning the pitch angle, with the optimal pitch angle range being approximately $30^{\circ}$ to $45^{\circ}$. Therefore, the flagellum only needs to make minor adjustments to its morphology to adapt to different living environments, which is why its Young’s modulus is relatively high.